# Large Language Models Might Not Care What You Are Saying: Prompt Format Beats Descriptions

### v6

@@ Appendix B Attention Weights on Alpaca
Appendix C Dataset Details for Reasoning Tasks
@@
C.1 Datasets Fetched from Exclusive Source
@@
C.2 Datasets Fetched from The Big-bench
@@
Appendix D Prompts for Reasoning Tasks Used in this Work
@@
Figure 18-26 show the examples of ERR (w/ CoT) prompt for respective datasets. Some tasks contain Answer Choices. In order to save space, the blank lines between the options are replaced with spaces in those figures. Each Figure has grey text for reasoning, cyan text for the parts of ERR that are unique to Vanilla, and italic words in the cyan text representing random nouns. Therefore, deleting the grey text gives ERR (w/o CoT), keeping the grey text but deleting the cyan text gives Vanilla (w/ CoT), and deleting both the cyan and grey text gives Vanilla (w/o CoT). [^4] When changing ”w/ CoT” to ”w/o CoT”, you may also need to replace ”So the answer is …” with ”The answer is …” for syntactical reasons. The reasoning is generated by ChatGPT  [^5] https://chatgpt.com/. Note the ChatGPT is not the same as GPT-3.5 we used for experiments.
Appendix E Results of Llama2
@@
Results of Llama2-7B-chat-hf Touvron et al. (2023b) on the nine datasets are presented in Figure 27. While ERR outperforms Vanilla with Llama2 across most datasets, its performance on Logical Fallacy and Sports is notably poor. Llama2 almost always responds with confused emojis for Logical Fallacy questions and outputs questions like "plausible or implausible?" for Sports, leading to predominantly incorrect answers. Further investigation into these issues is left for future work.
Appendix F Computational Details
F.1 Hardware

### v7

@@ Appendix C Attention Weights on Alpaca
Appendix D Dataset Details for Reasoning Tasks
@@
D.1 Datasets Fetched from Exclusive Source
@@
D.2 Datasets Fetched from The Big-bench
@@
Appendix E Prompts for Reasoning Tasks Used in this Work
@@
Figure 14-22 show the examples of ERR (w/ CoT) prompt for respective datasets. Some tasks contain Answer Choices. In order to save space, the blank lines between the options are replaced with spaces in those figures. Each Figure has grey text for reasoning, cyan text for the parts of ERR that are unique to Vanilla, and italic words in the cyan text representing random nouns. Therefore, deleting the grey text gives ERR (w/o CoT), keeping the grey text but deleting the cyan text gives Vanilla (w/ CoT), and deleting both the cyan and grey text gives Vanilla (w/o CoT). [^5] When changing ”w/ CoT” to ”w/o CoT”, you may also need to replace ”So the answer is …” with ”The answer is …” for syntactical reasons. The reasoning is generated by ChatGPT  [^6] https://chatgpt.com/. Note the ChatGPT is not the same as GPT-3.5 we used for experiments.
Appendix F Results of Llama2
@@
Results of Llama2-7B-chat-hf Llama Team (2023) on the nine datasets are presented in Figure 23. While ERR outperforms Vanilla with Llama2 across most datasets, its performance on Logical Fallacy and Sports is notably poor. Llama2 almost always responds with confused emojis for Logical Fallacy questions and outputs questions like "plausible or implausible?" for Sports, leading to predominantly incorrect answers. Further investigation into these issues is left for future work.
Appendix G Computational Details
G.1 Hardware